Given: C = 260 + 0.8Y I = 320 G = 300

a) Determine the equilibrium level of income using expenditure and injection-leakage approach

b) Calculate the value of mulitiplier for this economy
1
Expert's answer
2021-07-12T11:44:16-0400

PART A

Equilibrium Level"-C+(1-b)Y=G"

-260+(1-0.8)Y= 300

0.2Y=320+260

0.2Y=580

(0.2/0.2)Y=(580/0.2)

Y=2900

PART B

Multiplier value ;

"K=1\/(1-MPC)"

K= 1/(1-0.8)

K=1/0.2

K=5